Find the surface area of the part of the surface 2y+ 4z-x^2= 5 that lies above the triangle with vertices (0,0),(2,0), and (2,4).
1
Expert's answer
2020-11-04T11:23:30-0500

"\displaystyle

\textsf{If surface}\, S\, \textsf{has equation of the form}\\
z = f(x, y), \, \textsf{where}\, (x, y) \in R\, \textsf{and}\, f\, \textsf{has continuous}\\ \textsf{partial derivatives, the surface area}\\\textsf{is given by}\\


A = \iint_R \sqrt{1 + \left(\frac{\partial z}{\partial x}\right)^2 + \left(\frac{\partial z}{\partial y}\right)^2} \, \mathrm{d}A\\

\begin{aligned}
2y + 4z - x^2 &= 5\\
4z &=5 + x^2 - 2y\\
z &= \frac{5}{4} + \frac{x^2}{4} - \frac{y}{2}
\end{aligned}\\

\frac{\partial z}{\partial x} = \frac{x}{2}, \, \frac{\partial z}{\partial y} = -\frac{1}{2}\\

\therefore A = \iint_R \sqrt{1 + \frac{x^2}{4} + \frac{1}{4}} \, \mathrm{d}A\\


A = \frac{1}{2}\iint_R \sqrt{5 + x^2} \, \mathrm{d}A\\


\textsf{Equation of line AC is}\, y = 2x,\\
x \, \textsf{is integrated from}\, 0 \, \textsf{to}\, 2. \\

\begin{aligned}
\therefore A &= \frac{1}{2}\int_0^2 \int_0^{2x} \sqrt{5 + x^2} \,\mathrm{d}y\,\mathrm{d}x\\
&= \frac{1}{2}\int_0^2 y\sqrt{5 + x^2} \,\vert_{y=0}^{y=2x}\mathrm{d}x\\
&= \frac{1}{2}\int_0^2 2x\sqrt{5 + x^2} \mathrm{d}x\\
&=\int_0^2 x\sqrt{5 + x^2} \mathrm{d}x \\
&=\frac{1}{2}\int_0^2 \sqrt{5 + x^2} \,\mathrm{d}(5 + x^2)\\
&=\frac{1}{2}\cdot \frac{2}{3} (5 + x^2)^{\frac{3}{2}}\biggr\vert_0^2\\
&=\frac{(5 + 4)^{\frac{3}{2}} - 5^{\frac{3}{2}}}{3} \\
&= \frac{27 - 5\sqrt{5}}{3}
\end{aligned}"